OpenStudy (anonymous):

Solve for s. s divided by nine minus twenty-three is less than or equal to thirty-seven 2 7 one hundred twenty-six five hundred forty

OpenStudy (skullpatrol):

(s/9) -23 <= 37

OpenStudy (anonymous):

five hundred forty

OpenStudy (anonymous):

\[9\left(\frac{s}{9}-23\right)\leq (37)9 \]\[-207+s\leq 333 \]\[s\leq 540 \]

OpenStudy (skullpatrol):